OpenStudy (anonymous):

Fan and Medal! The graph of the sine curve below is of electromagnetic energy that represents orange light:

OpenStudy (anonymous):

sine graph with points at 0, 0 and 152 and 5 tenths, 1 and 305, 0 and 457 and 5 tenths, negative 1 and 610, 0. What function accurately represents the sine curve for orange light?

OpenStudy (anonymous):

@ImaGuest @ImNotGoodAtMath @geerky42 @Hero @Here_to_Help15

OpenStudy (anonymous):

Like I said before, I'm not good at math. Please stop tagging me

OpenStudy (anonymous):

please help @ShailKumar

OpenStudy (anonymous):

First of all, it should not be EM energy graph as energy is negative. It is an EM field graph. Let the function be y = A*sin (k*x) Put the given points to find A, k.

OpenStudy (anonymous):

does it matter what points i use?

OpenStudy (anonymous):

You can use any point

OpenStudy (anonymous):

a. f(x) = sin pi/610x b. f(x) = sin pi/305x c. f(x) = sin 610πx d. f(x) = sin 305πx

OpenStudy (anonymous):

y = 1sin(k * x) k = pi/610??? would the answer be A?

OpenStudy (anonymous):

Hang on..

OpenStudy (anonymous):

You have the points as (0,0), (152.5,1),(305,0),(457.5,-1) and (610,0) Put these points

OpenStudy (anonymous):

For instance 1 = A*sin (k *152.5)

OpenStudy (anonymous):

-1 = A*sin (k *457.5)

OpenStudy (anonymous):

Divide these two equation to solve for k.

OpenStudy (anonymous):

NOt to be rude but this is the last question i need help on and is there any way i could just get the answer.

OpenStudy (anonymous):

OK. What is the maximum value on y axis in the graph ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Great.. that is A.

OpenStudy (anonymous):

What is the wavelength ?

OpenStudy (anonymous):

there are no measurements on the x-axis so im not sure

OpenStudy (anonymous):

There is..! It is 610 nm

OpenStudy (anonymous):

\(k = 2\pi / wavelength\) Now k ?

OpenStudy (anonymous):

a. f(x) = sin pi/610x b. f(x) = sin pi/305x c. f(x) = sin 610πx d. f(x) = sin 305πx

OpenStudy (anonymous):

\(k = \pi / 305\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

What is the correct answer ?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Great! You are awesome ! :)

OpenStudy (anonymous):

no u r! tthanks!!